OpenStudy (fibonaccichick666):

Just a hint please. Solve: \[y'=\frac{t^2+y^2}{ty}\]

OpenStudy (fibonaccichick666):

I don't think I can use separable variables. I cant see integrating factor...

OpenStudy (matlee):

whats your question? sorry

OpenStudy (matlee):

o

OpenStudy (matlee):

\[y'=\frac{t^2+y^2}{ty}\]

OpenStudy (fibonaccichick666):

Solving the question stated

OpenStudy (fibonaccichick666):

but I only want a hint

OpenStudy (anonymous):

are u trying to fidn the anti derivative?

OpenStudy (fibonaccichick666):

no, it's an ODE

OpenStudy (fibonaccichick666):

I'm solving for y.

OpenStudy (amistre64):

split the fraction ...

OpenStudy (matlee):

i just did it i dont understand it either

OpenStudy (fibonaccichick666):

I did that, but it left me with nothing helpful

OpenStudy (fibonaccichick666):

\[y'=\frac{t}{y}+\frac{y}{t}\]

OpenStudy (amistre64):

try letting y/t = v, therefore vt=y

OpenStudy (fibonaccichick666):

k give me a minute

OpenStudy (fibonaccichick666):

so \(v'=v^{-1}+v\)?

OpenStudy (fibonaccichick666):

so then \[v'=\frac{1+v^2}{v}\]

OpenStudy (amistre64):

let me chk y' = v't + v v't + v = v^(-1) + v

OpenStudy (fibonaccichick666):

then I can use separable

OpenStudy (fibonaccichick666):

brilliant!!! Thank you!!!!

OpenStudy (amistre64):

\[v't+v=v^{-1}+v\] \[v't=v^{-1}\] \[v ~dv=\frac1t dt\] is what im getting, or do we see an error in it?

OpenStudy (fibonaccichick666):

I didn't get that

OpenStudy (amistre64):

if y = vy what is y'?

OpenStudy (fibonaccichick666):

so y=vt y'=v'

OpenStudy (amistre64):

vt is a prduct

OpenStudy (fibonaccichick666):

yep, but we are differentiating wrt t no?

OpenStudy (amistre64):

yes, but its still a product of 2 function related by t

OpenStudy (fibonaccichick666):

awe phooey

OpenStudy (amistre64):

x f(x) how wold we diffrentiate this?

OpenStudy (fibonaccichick666):

give me a minute to fix it

OpenStudy (fibonaccichick666):

now I get the same as you

OpenStudy (fibonaccichick666):

\(vdv=\frac{1}{t} dt\)

OpenStudy (fibonaccichick666):

I got it from here, will you check my final result?

OpenStudy (amistre64):

sure, but i usually mess it up from here :)

OpenStudy (anonymous):

You can check to see if the ODE is homogeneous using the following sort of procedure. \(y'=f(y,t)\) is homogeneous if you can write \(f(\lambda y,\lambda t)=\lambda^k f(y,t)\) for some real \(k\). If you can establish this, the substitution amistre suggested will always work.

OpenStudy (fibonaccichick666):

ok so I got,\[ y=t\sqrt{2ln|t|+c_1}\]

OpenStudy (amistre64):

v^2 = ln|t| + c v = sqrt(ln|t| + c) y/t = sqrt(ln|t| + c) y = t sqrt(ln|t| + c) kinda wonder where you got that 2ln(t) from

OpenStudy (fibonaccichick666):

integral of v is 1/2 v^2

OpenStudy (amistre64):

thnx sith, these things work by way of repressed memorys for me :)

OpenStudy (amistre64):

doh .... and 2ln|t| = ln(t^2) if you feel like it

OpenStudy (fibonaccichick666):

yea, but it's unnecessary since i'm not doing any e^ln stuff

OpenStudy (fibonaccichick666):

Thank you! I appreciate your help!

OpenStudy (amistre64):

your welcome ... srry i forgot how to integrate v :) knew id mess it up on this end somehow

OpenStudy (fibonaccichick666):

haha, you got the important part though!